OpenStudy (anonymous):

Find the common ratio for the following geometric sequence. 2, 1, 0.5, 0.25 Find the next term in the sequence. 128, 64, 32, Write the related series for the finite sequence 1.5, 2.7, 3.9, . . ., 8.7? Then find the sum. The given series has six terms. What is the sum of the terms of the series? 5 + 8 + 11 + . . . + 20

OpenStudy (anonymous):

c.r.=an/an-1=1/2=0.5 (for G.P) c.d for A.P=an-an-1